If x-5=2, what is the value of 3x-5?

A fish tank contains 25 trout,20 bass, 15 perch, and no other kinds of fish. If one of theses fish is to be chosen at random, what is the possibility that the fish chose will be a trout?
if (x-3)(x+3)=a, then ( 2x-6)(x+3)=?
Which of the following could be represented by the expression 7-n/2?

Found 2 solutions by nyc_function, solver91311:
Answer by nyc_function(2741)

BIG TIP: Post one question at a time or else no will reply.

If x-5=2, what is the value of 3x-5?

x - 5 = 2

x = 2 + 5

x = 7
====================================

Let x = 7

3x - 5 =

3(7) - 5 =

21 - 5 = 16

Post the other two questions separately.

Answer by solver91311(24713)



One question per post.



But



by factoring a 2 out of the first binomial. But since





John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism